8. (a) Write the dual of the following LPP: (5)

Minimize 16 1 9 2 21 3 Z = x + x + x

Subject to the constraints

x1 + x2 + x3 = 16

2x1 + x2 + x3 ≥12

x1

, x2 ≥ 0

3

x -unrestricted.


1
Expert's answer
2021-05-07T11:00:02-0400

Max "z=7x_1+12x_2+4x_3"

   subject to-

       "x_1+x_2+x_3=16\\

 2x_1+x_2+x_3\ge 12\\

 x_1,x_2,x_3\ge 0"

Dual is-

     Objective function- min  "10y_1+8y_2"


   constraints would be-

   "y_1+2y_1\ge 7\\2y_1-y_2\ge 12\\y_1+3y_2\ge 4\\ y_1,y_2\ge 0"


This is the dual of the Primal Problem.